On the blueprint, the scale indicates that 2 cm represents 5 ft. What are the actual dimensions of a room that is 4.8 cm by 6 cm on the blueprint?
Found 2 solutions by stanbon, josmiceli:
Answer by stanbon(48558)

(4.8/2)(5 ft)= 2.4 * 5 = 12 ft.
(6/2)(5ft)= 3(5)=15 ft.
Cheers,
Stan H.

Answer by josmiceli(6786)

2 cm represents 5 ft.
5ft/2cm = 2.5 ft/1cm
square top and bottom of this ratio including the units
6.25 square ft/1 square cm
(6.25 square ft/1 square cm)(4.8 cm x 6 cm) =
(6.25 square ft/1 square cm)(28.8 square cm)=
180 square ft